# Smooth orbit equivalence rigidity for dissipative geodesic flows

### v1

@@ 2.2. Complex geometry
The Teichmüller space of $M$, denoted by $\mathcal{T}(M)$, is the space of complex structures on $M$ modulo the equivalence relation that $J\sim\tilde{J}$ if and only if there exists a diffeomorphism $\psi:M\to M$, isotopic to the identity, such that $\psi^{*}\tilde{J}=J$. We will denote such an equivalence class of complex structures by $[J]$.
The mapping class group $\text{MCG}(M)$ is defined as the quotient of orientation preserving diffeomorphisms on $M$ modulo isotopy. They act on $\mathcal{T}(M)$ by pullback, and the quotient space $\mathcal{M}(M):=\mathcal{T}(M)/\text{MCG}(M)$ is the moduli space of complex structures on $M$. See [FM11] for a thorough introduction.
Each complex structure $J$ determines a canonical line bundle $\kappa:=T^{\ast}_{1,0}M$ on $M$. We will denote by $H^{0}_{J}(M,\kappa^{\otimes k})$ the space of $J$-holomorphic sections of the $k$-th tensor power of the canonical line bundle $\kappa$. Locally, its elements have the form $w(z)dz^{k}$ for $k\geq 0$ and $w(z)d\bar{z}^{-k}$ for $k<0$.
@@ 2.2.2. Fiberwise holomorphic distributions
Each subspace $\Omega_{k}$ of Fourier modes can be identified with $\mathcal{C}^{\infty}(M,\kappa^{\otimes k})$, the set of smooth sections of the bundle $\kappa^{\otimes k}$. Indeed, we have a $\mathbb{C}$-linear isomorphism
@@
given by restriction to $SM$, i.e., in local coordinates (for $k\geq 0$),
@@
This is a generalization of the map $\pi_{1}^{*}$ which we have already encountered in (1.2) to identify smooth $1$-forms with $\Omega_{-1}\oplus\Omega_{1}$. Note that the definition of $\pi_{k}^{\ast}$ depends on the choice of the metric $g$. We denote by $\pi_{k\ast}$ its $L^{2}$-adjoint. In local coordinates, we have
$(\pi_{k\ast}u)(x)=\left(\int_{\pi^{-1}(x)}u(x,\cdot)\right)dz^{k}.$
Once we extend the operators to distributions by duality, the projection onto the $k$-th Fourier mode is simply given by $(2\pi)^{-1}\pi^{\ast}_{k}\pi_{k\ast}$ acting on $\mathcal{D}^{\prime}(SM)$.
Under this identification, we can essentially think of the raising/lowering operators $\eta_{\pm}$ as $\bar{\partial}$ and $\partial$ operators thanks to the following result (see [PSU14, Lemma 2.1] and the ensuing discussion).
@@
For $k\geq 0$, the following diagram commutes:
@@
We also introduce the following terminology:
@@
A distribution $u\in\mathcal{D}^{\prime}(SM)$ is said to be fiberwise holomorphic if $u_{k}=0$ for all $k<0$.
Equivalently, if we define the Szegö projectors $S_{\pm}:\mathcal{D}^{\prime}(SM)\to\mathcal{D}^{\prime}(SM)$ by
$S_{+}u=\sum_{k\geq 0}u_{k},\qquad S_{-}u=\sum_{k\leq 0}u_{k},$
then a distribution $u$ is fiberwise holomorphic if and only if $S_{+}u=u$. The projectors satisfy the commutation relations
@@
We will be interested in the family of fiberwise holomorphic distributions that satisfy the transport equation:
$\mathcal{D}^{\prime}_{\text{tr},+}(SM):=\{u\in\mathcal{D}^{\prime}(SM)\mid(F+V(\lambda))u=0,\,S_{+}u=u\}.$ (2.6)
@@ 2.2.3. Torelli’s theorem
The following form of Torelli’s theorem tells us that period matrices capture a lot of the information about the complex structure.
@@
Assume that $M$ has genus $\geq 2$. If $\Pi(J)=\Pi(\tilde{J})$, then there exists an orientation-preserving diffeomorphism $\psi:M\to M$ such that $\psi^{*}\tilde{J}=J$.
@@ 2.3. Hyperbolic dynamics
We now further assume that the flow of the generalized thermostat $(M,g,\lambda)$ is Anosov (or uniformly hyperbolic).
@@ 2.3.1. Definition
Recall that the Anosov property means that there exists a flow-invariant continuous splitting
$T(SM)=\mathbb{R}F\oplus E^{s}\oplus E^{u}$
@@
$\|d\varphi_{t}|_{E^{s}}\|\leq C\rho^{t},\qquad\|d\varphi_{-t}|_{E^{u}}\|\leq C\rho^{t}.$ (2.8)
In the geodesic case, the contact form $\alpha$ is preserved, so $\ker\alpha=\mathbb{H}\oplus\mathbb{V}=E^{s}\oplus E^{u}$. It is then known that $E^{s}\cap\mathbb{V}=\{0\}=E^{u}\cap\mathbb{V}$. For a generalized thermostat, we instead know by [DP07, Lemma 4.1] that
$(\mathbb{R}F\oplus E^{s})\cap\mathbb{V}=\{0\}=(\mathbb{R}F\oplus E^{u})\cap\mathbb{V}.$ (2.9)
Here $\mathbb{R}F\oplus E^{s/u}$ are the weak stable and unstable bundles. This implies that there exist $r^{s/u}\in\mathcal{C}^{0}(SM,\mathbb{R})$ such that
$Y^{s}:=H+r^{s}V\in\mathbb{R}F\oplus E^{s},\qquad Y^{u}:=H+r^{u}V\in\mathbb{R}F\oplus E^{u}.$ (2.10)
In fact, the weak stable and unstable bundles are $\mathcal{C}^{1}$ (see [Has94, Corollary 1.8]), so the functions $r^{s/u}$ are also $\mathcal{C}^{1}$ (and smooth along the flow since each bundle $\mathbb{R}F\oplus E^{s/u}$ is $\varphi_{t}$-invariant). The Anosov property implies that $r^{s}\neq r^{u}$ everywhere. One may in fact show that $r^{s}<r^{u}$, so the basis $\{F,Y^{s},Y^{u}\}$ is positively-oriented.
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. Then, the functions $r^{s/u}\in\mathcal{C}^{1}(SM,\mathbb{R})$ uniquely characterized by (2.10) satisfy $r^{s}<r^{u}$.
@@
Since $r^{s}\neq r^{u}$ everywhere, it suffices to show the inequality at a single point. By compactness, we can pick $(x,v)\in SM$ such that $V(\lambda)(x,v)=0$. Let us define
$\xi(t):=d\varphi_{-t}(V_{\varphi_{t}(x,v)}).$
@@
$\dot{\xi}(0)=-H_{(x,v)}.$
Since $r^{s}(x,v)\neq r^{u}(x,v)$, there exists a unique constant $c\in\mathbb{R}$ such that $V_{(x,v)}+cX_{(x,v)}$ belongs to $E^{s}\oplus E^{u}$. Therefore, given that $E^{s}$ and $E^{u}$ are uniformly attracting and repelling sets on $E^{s}\oplus E^{u}$ respectively, we must have $r^{s}(x,v)<r^{u}(x,v)$ at this point.
@@
Note that, when $\lambda=0$ and $K_{g}<0$, i.e., in the geodesic case with negative curvature, we have the stronger statement $r^{s}<0<r^{u}$ because $[X,H]=K_{g}V$.
The dual bundles are defined by
$(\mathbb{R}F)^{*}(E^{s}\oplus E^{u})=0=(E^{s})^{\ast}(\mathbb{R}F\oplus E^{s})=(E^{u})^{\ast}(\mathbb{R}F\oplus E^{u}).$
One can check we have similar estimates to (2.8) for $(E^{s})^{*}$ and $(E^{u})^{*}$, with $d\varphi_{t}$ replaced by $d\varphi_{t}^{-T}$ (inverse transpose). Translated to the setting of the cotangent bundle, property (2.9) then becomes
$(E^{s})^{\ast}\cap\mathbb{H}^{*}=\{0\}=(E^{u})^{\ast}\cap\mathbb{H}^{\ast}.$ (2.11)
@@
Figure 1. The relevant subbundles in tangent and cotangent spaces.
@@ 2.3.2. Tensor tomography
We say that a thermostat $(M,g,\lambda)$ satisfies the attenuated tensor tomography problem of order $n$ if having $(F+V(\lambda))u=f$ with $f,u\in\mathcal{C}^{\infty}(SM)$ and $f$ of Fourier degree $n\geq 0$ implies that $u$ is of Fourier degree $\max(n-1,0)$.
The term ‘attenuated’ refers to the presence of the divergence $V(\lambda)$ in the transport equation. Note that such a term appears for Gaussian thermostats and generalized thermostats of higher Fourier degree, but not for magnetic or geodesic flows.
The fact that geodesic flows satisfy the (attenuated) tensor tomography problem was first proved in negative curvature in [GK80] for $n\geq 0$ and then generalized to the Anosov case in [DS03] for $n\leq 1$, [PSU14] for $n\leq 2$, and [Gui17] for $n\geq 2$. It was also shown in [DP05] that Anosov magnetic flows satisfy the (attenuated) tensor tomography problem of order $n\leq 1$.
For generalized thermostats of higher Fourier degree, the non-attenuated and attenuated versions of the tensor tomography problem are different. In [DP07], it was proved that Gaussian thermostats (potentially mixed with a magnetic component) satisfy the non-attenuated tensor tomography problem of order $n\leq 1$. We instead need:
@@
This result is a consequence of the work in [AR21]. Their argument heavily relies on the negative thermostat curvature assumption. In particular, most of the heavy lifting is done by [AR21, Theorem 3.1], where the Carleman estimates for Gaussian thermostats with negative curvature are established (akin to the work in [PS23]):
@@
Let $(M,g,\theta)$ be a Gaussian thermostat with $\mathbb{K}<0$. Suppose $f\in\mathcal{C}^{\infty}(SM)$ has finite Fourier degree and $u\in\mathcal{C}^{\infty}(SM)$ satisfies $(F+V(\lambda))u=f$. Then $u$ also has finite Fourier degree.
@@
We follow the argument from [AR21, Theorem 5.1]. Let $m^{\prime}\geq 0$ be the Fourier degree of $f$. Since $(F+V(\lambda))u=f$, we obtain
$(Fu)_{k}=-i\lambda_{1}u_{k-1}+i\lambda_{-1}u_{k+1}\quad\text{ for all }|k|\geq m^{\prime}+1.$
@@
$\|(Fu)_{k}\|^{2}\leq C(\|u_{k-1}\|^{2}+\|u_{k+1}\|^{2})\quad\text{ for all }|k|\geq m^{\prime}+1.$
@@
However, we have $01-eC/(\kappa s)>$ 0 by design, so $u_{k}=0$ for all $|k|\geq m$.
@@
By Proposition 2.14, we know that $u$ is of finite Fourier degree. Suppose, for the sake of contradiction, that $u$ is of degree $k\geq 1$. Then, using the equation $(F+V(\lambda))u=f$, we have
@@
Finally, the proofs of our theorems rely on the possibility of lifting arbitrary holomorphic $1$-forms to solutions of the transport equation. As explained in Appendix A, where we have relegated most of the work on this front, this is again related to the injectivity of the X-ray transform for thermostats.
@@
Let $(M,g,\lambda)$, with $\lambda$ of Fourier degree $\leq 1$, be an Anosov thermostat. For any holomorphic (resp. anti-holomorphic) $1$-form $\tau$ on $M$, there exists $u\in H^{-1}(SM)$ with $u_{k}=0$ for all $k\leq 0$ (resp. $k\geq 0$) such that $(F+V(\lambda))u=0$ and $u_{1}=\pi_{1}^{*}\tau$ (resp. $u_{-1}=\pi_{1}^{*}\tau$).
@@
Let us treat the case where the $1$-form $\tau$ is holomorphic. The anti-holomorphic case is completely analogous. Using Lemma 2.5, we know that $\pi_{1}^{*}\tau\in\Omega_{1}$ is in the kernel of $\eta_{-}$. We can hence apply Theorem A.5, which tells us that there exists $v\in H^{-1}(SM)$ with $(F+V(\lambda))v=0$ such that $v_{1}=\pi_{1}^{*}\tau$ and $v_{-1}=v_{0}=0$. We project this distribution onto its positive Fourier components to get $u:=S_{+}v=\sum_{k\geq 1}v_{k}$. For all $k\in\mathbb{Z}$, we then have
$((F+V(\lambda))u)_{k}=\eta_{+}u_{k-1}+\eta_{-}u_{k+1}+ik(\lambda_{1}u_{k-1}+\lambda_{0}u_{k}+\lambda_{-1}u_{k+1})=0,$
which entails that $(F+V(\lambda))u=0$. ∎
@@
Suppose $\lambda=\lambda_{m}+\lambda_{-m}$ where $m\geq 2$ and $\eta_{-}\lambda_{-m}=0$. Since $\lambda_{-m}$ has isolated zeroes, there exists $a\in\Omega_{1}$ with $\eta_{-}a=0$ and $\lambda_{-m}a\neq 0$. Then, there is no $u\in H^{-1}(SM)$ with $u_{k}=0$ for all $k\leq 0$ such that $(F+V(\lambda))u=0$ and $u_{1}=a$.
@@
Suppose such a distribution $u$ exists. For any $k\in\mathbb{Z}$, we must have
$0=((F+V(\lambda))u)_{k}=\eta_{+}u_{k-1}+\eta_{-}u_{k+1}+ik(\lambda_{m}u_{k-m}+\lambda_{-m}u_{k+m}).$
Therefore, applying this to $k=-m+1$, we get $\lambda_{-m}a=\lambda_{-m}u_{1}=0$, a contradiction.∎

### v2

@@ 2.2. Complex geometry
The Teichmüller space of $M$, denoted by $\mathcal{T}(M)$, is the space of complex structures on $M$ modulo the equivalence relation that $J\sim\widetilde{J}$ if and only if there exists a diffeomorphism $\psi:M\to M$ isotopic to the identity such that $\psi^{*}\widetilde{J}=J$. We will denote such an equivalence class of complex structures by $[J]$.
The mapping class group $\text{MCG}(M)$ is defined as the quotient of orientation-preserving diffeomorphisms on $M$ modulo isotopy. They act on $\mathcal{T}(M)$ by pullback, and the quotient space $\mathcal{M}(M)\coloneqq\mathcal{T}(M)/\text{MCG}(M)$ is the moduli space of complex structures on $M$. See [FM12] for a thorough introduction.
Each complex structure $J$ determines a canonical line bundle $\kappa:=T^{\ast}_{1,0}M$ and its conjugate $\overline{\kappa}:=T^{\ast}_{0,1}M$ over $M$. Locally, the sections of $\kappa$ have the form $w(z)\,dz$ while the sections of $\overline{\kappa}$ have the form $w(z)\,d\overline{z}$. The dual of $\kappa$ is called the anti-canonical line bundle and we identify it with $\overline{\kappa}$ by using the Hermitian inner product on $\kappa$ induced by the Riemannian metric on $M$. We then denote this bundle by $\kappa^{-1}$ so that the tensor powers $\kappa^{\otimes k}$ make sense for all $k\in\mathbb{Z}$.
We will denote by $H^{0}_{J}(M,\kappa^{\otimes k})$ the space of $J$-holomorphic sections of the $k$th tensor power of the canonical line bundle $\kappa$. Locally, its elements have the form $w(z)\,dz^{k}$ for $k\geq 0$ and $w(z)\,d\bar{z}^{-k}$ for $k<0$.
@@ 2.2.2. Fiberwise holomorphic distributions
Each subspace $\Omega_{k}$ of Fourier modes can be identified with $\mathcal{C}^{\infty}(M,\kappa^{\otimes k})$, the set of smooth sections of the bundle $\kappa^{\otimes k}$ [PSU23, Lemma 6.1.19]. Indeed, we have a $\mathbb{C}$-linear isomorphism
@@
given by restriction to $SM$, that is, locally, for $k\geq 0$,
@@
Note that the definition of $\pi_{k}^{\ast}$ depends on the choice of the metric $g$. We denote by $\pi_{k\ast}$ its $L^{2}$ adjoint. Locally, for $k\geq 0$, we have
$(\pi_{k\ast}u)(x)=\left(\int_{S_{x}M}u(x,\cdot)\psi\right)\,dz^{k}.$
Once we extend the operators to distributions by duality, the projection onto the $k$th Fourier mode is simply given by $(2\pi)^{-1}\pi^{\ast}_{k}\pi_{k\ast}$ acting on $\mathcal{D}^{\prime}(SM)$.
Under this identification, we can essentially think of the raising/lowering operators $\eta_{\pm}$ as $\partial$ and $\bar{\partial}$ operators thanks to the following result (see [PSU23, pp. 153–154]).
@@
For $k\geq 0$, the following diagram commutes.
@@
We also introduce the following terminology.
@@
A distribution $u\in\mathcal{D}^{\prime}(SM)$ is said to be fibrewise holomorphic if $u_{k}=0$ for all $k<0$.
Equivalently, if we define the Szegő projectors $S_{\pm}:\mathcal{D}^{\prime}(SM)\to\mathcal{D}^{\prime}(SM)$ by
$S_{+}u\coloneqq\sum_{k\geq 0}u_{k},\qquad S_{-}u\coloneqq\sum_{k\leq 0}u_{k},$
then a distribution $u$ is fibrewise holomorphic if and only if $S_{+}u=u$. The projectors satisfy the commutation relations
@@
We will be interested in the family of fibrewise holomorphic distributions that satisfy the transport equation:
$\mathcal{D}^{\prime}_{\text{tr},+}(SM)\coloneqq\left\{u\in\mathcal{D}^{\prime}(SM)\mid(F+V\lambda)u=0,\,S_{+}u=u\right\}.$ (2.6)
@@ 2.2.3. Torelli’s theorem
The following form of Torelli’s theorem tells us that period matrices capture a lot of information about the complex structure.
@@
Assume that $M$ has genus $\geq 2$. If $\Pi(J)=\Pi(\widetilde{J})$, then there exists an orientation-preserving diffeomorphism $\psi:M\to M$ such that $\psi^{*}\widetilde{J}=J$.
@@ 2.3. Hyperbolic dynamics
We now further assume that the flow of the thermostat $(M,g,\lambda)$ is Anosov (or uniformly hyperbolic).
@@ 2.3.1. Definition
Recall that the Anosov property means that there exist a flow-invariant continuous splitting
$T(SM)=\mathbb{R}F\oplus E_{s}\oplus E_{u}$
@@
$\left\|d\varphi_{t}|_{E_{s}}\right\|\leq C\rho^{t},\qquad\left\|d\varphi_{-t}|_{E_{u}}\right\|\leq C\rho^{t}.$ (2.8)
In the geodesic case, the contact form $\alpha$ is preserved, so $\ker\alpha=\mathbb{H}\oplus\mathbb{V}=E_{s}\oplus E_{u}$. It is then known that $E_{s}\cap\mathbb{V}=\{0\}=E_{u}\cap\mathbb{V}$. For a thermostat, we instead know by [DP07, Lemma 4.1] that
$(\mathbb{R}F\oplus E_{s})\cap\mathbb{V}=\{0\}=(\mathbb{R}F\oplus E_{u})\cap\mathbb{V}.$ (2.9)
Here, $\mathbb{R}F\oplus E_{s/u}$ are the weak stable and unstable subbundles. This implies that there exist $r^{s},r^{u}\in\mathcal{C}^{0}(SM,\mathbb{R})$ such that
$Y^{s}\coloneqq H+r^{s}V\in\mathbb{R}F\oplus E_{s},\qquad Y^{u}\coloneqq H+r^{u}V\in\mathbb{R}F\oplus E_{u}.$ (2.10)
In fact, the weak stable and unstable subbundles are $\mathcal{C}^{1}$ (see [Has94, Corollary 1.8]), so the functions $r^{s}$ and $r^{u}$ are also $\mathcal{C}^{1}$ (and smooth along the flow since each subbundle $\mathbb{R}F\oplus E_{s/u}$ is flow-invariant). The Anosov property implies that $r^{s}\neq r^{u}$ everywhere. One may in fact show that $r^{s}<r^{u}$, so the global frame $(F,Y^{s},Y^{u})$ is positively oriented. See Figure 1.
@@
Let $(M,g,\lambda)$ be an Anosov thermostat. Then, the differentiable functions $r^{s},r^{u}\in\mathcal{C}^{1}(SM,\mathbb{R})$ uniquely characterized by equation (2.10) satisfy $r^{s}<r^{u}$.
@@
Since $r^{s}\neq r^{u}$ everywhere, it suffices to show the inequality at a single point. By compactness, we can pick $(x,v)\in SM$ such that $(V\lambda)(x,v)=0$. Let us define
$\zeta(t)\coloneqq d\varphi_{-t}(V(\varphi_{t}(x,v))).$
@@
$\dot{\zeta}(0)=-H(x,v).$
Since $r^{s}(x,v)\neq r^{u}(x,v)$, there is a unique constant $c\in\mathbb{R}$ such that $V(x,v)+cX(x,v)$ belongs to $E_{s}\oplus E_{u}$. Therefore, since $E_{s}$ and $E_{u}$ are uniformly repelling and attracting sets on $E_{s}\oplus E_{u}$, respectively, we must have $r^{s}(x,v)<r^{u}(x,v)$ at this point. ∎
@@
When $\lambda=0$ and $K_{g}<0$, that is, in the geodesic case with negative curvature, we have the stronger statement $r^{s}<0<r^{u}$ because $[X,H]=\pi^{*}(K_{g})V$.
The dual subbundles are defined by
$E_{s}\ast(\mathbb{R}F\oplus E_{s})=0=E_{u}^{\ast}(\mathbb{R}F\oplus E_{u}).$
One can check that $E_{s}^{*}$ and $E_{u}^{*}$ satisfy analogues of the estimates (2.8), with $d\varphi_{t}$ replaced by $(d\varphi_{t}^{\top})^{-1}$ (inverse of the transpose). Translated to the setting of the cotangent bundle, property (2.9) then becomes
$E_{s}^{\ast}\cap\mathbb{H}^{*}=\{0\}=E_{u}^{\ast}\cap\mathbb{H}^{\ast}.$ (2.11)
@@
Figure 1. The relevant subbundles of the tangent and cotangent bundles.
@@ 2.3.2. Tensor tomography
We say that a thermostat $(M,g,\lambda)$ satisfies the attenuated tensor tomography problem of order $n$ if having $(F+V\lambda)u=f$ with $f,u\in\mathcal{C}^{\infty}(SM)$ and $f$ of Fourier degree $n\geq 0$ implies that $u$ is of Fourier degree $\max(n-1,0)$.
The term ‘attenuated’ refers to the presence of the divergence $V\lambda$ in the transport equation. Such a term appears for Gaussian thermostats and thermostats of higher Fourier degree, but not for magnetic or geodesic flows.
The fact that geodesic flows satisfy the (attenuated) tensor tomography problem was first proved in negative curvature in [GK80] for $n\geq 0$, and then generalized to the Anosov case in [DS03] for $n\leq 1$, [PSU14] for $n\leq 2$ and [Gui17] for $n\geq 2$. It was also shown in [DP05] that Anosov magnetic flows satisfy the (attenuated) tensor tomography problem of order $n\leq 1$.
For thermostats of higher Fourier degree, the non-attenuated and attenuated versions of the tensor tomography problem are different. In [DP07], it was proved that Gaussian thermostats (potentially mixed with a magnetic component) satisfy the non-attenuated tensor tomography problem of order $n\leq 1$. We instead need the following theorem.
@@
This result is a consequence of the work in [AR21]. Their argument relies heavily on the negative thermostat curvature assumption. In particular, most of the heavy lifting is done by [AR21, Theorem 3.1], where the Carleman estimates for Gaussian thermostats with negative curvature are established (akin to the work in [PS23]).
@@
Let $(M,g,\theta)$ be a Gaussian thermostat with $\mathbb{K}<0$. Suppose that $f\in\mathcal{C}^{\infty}(SM)$ has finite Fourier degree and $u\in\mathcal{C}^{\infty}(SM)$ satisfies $(F+V\lambda)u=f$. Then, the function $u$ also has finite Fourier degree.
@@
We follow the argument from [AR21, Theorem 5.1]. Let $m^{\prime}\geq 0$ be the Fourier degree of $f$. Since $(F+V\lambda)u=f$, we obtain
$(Fu)_{k}=-i\lambda_{1}u_{k-1}+i\lambda_{-1}u_{k+1}\qquad\text{for all }|k|\geq m^{\prime}+1.$
@@
$\|(Fu)_{k}\|^{2}\leq C(\|u_{k-1}\|^{2}+\|u_{k+1}\|^{2})\qquad\text{for all }|k|\geq m^{\prime}+1.$
@@
However, we have $01-eC/(\kappa s)>$ 0 by design, so $u_{k}=0$ for all $|k|\geq m$. ∎
@@
By Proposition 2.14, we know that $u$ is of finite Fourier degree. Suppose, for the sake of contradiction, that $u$ is of degree $k\geq 1$. Then, using the equation $(F+V\lambda)u=f$, we have
@@
Finally, the proofs of our results rely on the possibility of lifting arbitrary holomorphic $1$-forms to solutions of the transport equation. As explained in Appendix A, where we have relegated most of the work on this front, this is again related to the injectivity of the X-ray transform for thermostats.
@@
Let $(M,g,\lambda)$, with $\lambda$ of Fourier degree $\leq 1$, be an Anosov thermostat. For any holomorphic (respectively anti-holomorphic) $1$-form $\tau$ on $M$, there exists a distribution $u\in H^{-1}(SM)$ with $u_{k}=0$ for all $k\leq 0$ (respectively $k\geq 0$) such that $(F+V\lambda)u=0$ and $u_{1}=\pi_{1}^{*}\tau$ (respectively $u_{-1}=\pi_{-1}^{*}\tau$).
@@
Let us treat the case where the $1$-form $\tau$ is holomorphic. The anti-holomorphic case is completely analogous. Using Lemma 2.5, we know that $\pi_{1}^{*}\tau\in\Omega_{1}$ is in the kernel of $\eta_{-}$. We can hence apply Theorem A.5, which tells us that there exists $v\in H^{-1}(SM)$ with $v_{0}=0$ such that $(F+V\lambda)v=0$ and $v_{-1}+v_{1}=\pi_{1}^{*}\tau$. Since $\pi_{1}^{*}\tau\in\Omega_{1}$, we actually have $v_{-1}=0$. We project the distribution $v$ onto its positive Fourier modes to get $u\coloneqq S_{+}v=\sum_{k\geq 1}v_{k}$. For all $k\in\mathbb{Z}$, we then get
$((F+V\lambda)u)_{k}=\eta_{+}u_{k-1}+\eta_{-}u_{k+1}+ik(\lambda_{1}u_{k-1}+\lambda_{0}u_{k}+\lambda_{-1}u_{k+1})=0,$
which entails that $(F+V\lambda)u=0$. ∎
@@
Suppose that $\lambda=\lambda_{m}+\lambda_{-m}$, where $m\geq 2$ and $\eta_{-}\lambda_{-m}=0$. Since $\lambda_{-m}$ has isolated zeroes, there exists $a\in\Omega_{1}$ with $\eta_{-}a=0$ and $\lambda_{-m}a\neq 0$. Then, there is no $u\in H^{-1}(SM)$ with $u_{k}=0$ for all $k\leq 0$ such that $(F+V\lambda)u=0$ and $u_{1}=a$.
@@
Suppose that such a distribution $u$ exists. For any $k\in\mathbb{Z}$, we must have
$0=((F+V\lambda)u)_{k}=\eta_{+}u_{k-1}+\eta_{-}u_{k+1}+ik(\lambda_{m}u_{k-m}+\lambda_{-m}u_{k+m}).$
Therefore, applying this identity to $k=-m+1$, we get $\lambda_{-m}a=\lambda_{-m}u_{1}=0$, which is a contradiction.∎